# Geometric properties of the additional third-order transitions in the two-dimensional Potts model

### v2

@@ 3 Results and discussion
The following section discusses the results obtained from Swendsen-Wang spin-cluster simulations and cluster analysis to elucidate the system behavior associated with the additional transitions in the Potts model, as identified by microcanonical inflection-point analysis.
3.1 Critical transition
@@
Figure 3: The variation of $\langle P_{\infty}\rangle$ with temperature is shown. The black dashed line marks the precise solution of the system’s phase transition temperature $T_{c}$. As shown in panel (a) and (b), it is observed that at $T_{c}$, the $q=3,4$ Potts models exhibit characteristics of a continuous phase transition, indicating that the system undergoes a second-order phase transition for $q=3,4$. In contrast, in panel (c) and (d), for $q=6,8$, $\langle P_{\infty}\rangle$ shows a discontinuity at $T_{c}$, signifying that the system undergoes a first-order phase transition for $q=6,8$.
@@
The primary goal of this paper is to determine the locations of both the third-order dependent and independent transitions in the $q$-state Potts model (for $q\geq 3$) for finite system sizes. The system sizes studied in this work are based on square lattices with edge lengths $L=32,64,128,256$. As derived by Wu [39], the exact solution for the critical transition temperature $T_{c}$ of the $q$-state Potts model (for $q\geq 3$) is given by:
$T_{c}=\frac{1}{\ln(1+\sqrt{q})}$ (6)
@@
Therefore, for $q=3,4,6,$ and $8$, the corresponding critical transition temperatures approximate $T_{c}(3)\approx 0.995$, $T_{c}(4)\approx 0.910$, $T_{c}(6)\approx 0.808$, and $T_{c}(8)\approx 0.745$, respectively.
@@
After identifying all the clusters in the system, various types of clusters emerge, and it is inevitable that the system contains a largest spin cluster. The proportion of spins in this largest cluster relative to the total system size, $\langle n_{\text{largest}}/L^{2}\rangle$, defines the magnetization $\langle M\rangle$ in the ferromagnetic system [40]. By analyzing the behavior of $\langle M\rangle$, the order of the phase transition in the $q$-state Potts model (for $q\geq 3$) can be determined [46]. In Fig. 3 (a) and (b), it is observed that for $q=3$ and $q=4$, $\langle M\rangle$ changes continuously with temperature, which aligns with the results from Wu [39] and Baxter [11]. This confirms that the phase transitions in the $q$-state Potts model for $q=3$ and $q=4$ are critical, i.e., second-order phase transitions. For $q=6$ and $q=8$, $\langle M\rangle$ exhibits discontinuous behavior, consistent with the numerical simulation results reported in the existing literature [11, 13, 14, 15, 16, 17], indicating a first-order phase transition when $q=6$ and $q=8$.
@@
Figure 4: The variation of isolated spins (ISO) with temperature for $q=3,4,6,8$ is depicted in panels (a)-(d), respectively. The purple dashed line marks the temperature corresponding to the third-order independent transition, while the black dashed line indicates the precise solution of the system’s phase transition temperature $T_{c}$. As shown in the figure, the temperatures corresponding to the third-order independent transition are identified as $T_{\text{in}}(3)=0.964$, $T_{\text{in}}(4)=0.88$, $T_{\text{in}}(6)=0.788$, and $T_{\text{in}}(8)=0.734$.
3.2 Third-order independent transition in the ferromagnetic phase
@@
As shown in Fig. 4 (a)-(d), the number of isolated spins in the $q=3,4,6,8$ Potts models, respectively, varies as a function of temperature. It is observed that the number of isolated spins initially increases with temperature, reaching a peak at $T_{\text{in}}(3)\approx 0.964$ for the 3-state Potts model, $T_{\text{in}}(4)\approx 0.880$ for the 4-state Potts model, $T_{\text{in}}(6)\approx 0.788$ for the 6-state Potts model, and $T_{\text{in}}(8)\approx 0.734$ for the 8-state Potts model. In Fig. 4, comparing the temperature of the third-order independent transition, $T_{\text{ind}}$, with the phase transition temperature, $T_{c}$, at the same $q$, the differences are approximately $\Delta T(3)\approx 0.031$, $\Delta T(4)\approx 0.030$, $\Delta T(6)\approx 0.020$, and $\Delta T(8)\approx 0.011$. It is evident that for $q=3$ and $q=4$, the third-order independent transition occurs at a notable distance from the critical transition, with $\Delta T(3)\approx\Delta T(4)$. However, for $q=6$ and $q=8$, $\Delta T$ starts to decrease, and as $q$ increases, $\Delta T$ becomes smaller. Although the temperatures of the third-order independent transition and the system’s critical transition become closer, microcanonical inflection point analysis shows a re-entrant behavior of the inverse temperature as energy increases[47], meaning that a single temperature corresponds to two energy levels. Near $T=T_{c}$, the system exhibits a two-phase coexistence. Therefore, despite the proximity of these temperatures, there remains a distinct difference in the system’s states.
@@
Figure 5: A comparison of isolated spins (ISO) and single spins in the $q=3$ Potts model is provided on a lattice with $N=200\times 200$. In the figure, the largest cluster size(LCS) in the system is shaded in gray, while all other clusters containing more than two spins are depicted in white, regardless of their spin orientations. The red points represent spins with $s_{i}=0$, the green points represent $s_{i}=1$, and the blue points represent $s_{i}=2$.
@@
As shown in Fig.5, it becomes apparent that at low temperatures ($T<T_{c}$), the largest cluster (LCS) dominates the system, nearly occupying the entire space, with little variation in its size and configuration. Prior to the phase transition, the percolation cluster continues to dominate, oscillating between similar structures. Disrupting the LCS at low temperatures can be achieved by isolated spins with orientations different from those within the LCS. During the spin-flip process, the LCS has a probability $P=1-e^{-K}$ of being flipped, where $K=J/(k_{B}T)$, resulting in changes to the required isolated spin states. As temperature increases, more clusters form, leading to an increase in the number of isolated spins of various types. In the paramagnetic phase at high temperatures, the system becomes highly disordered, with more clusters appearing. Panels (a)-(d) show the variation of isolated spins with temperature: $T(a)=T(e)=0.8$, $T(b)=T(f)\approx T_{\text{in}}\approx 0.964$, $T(c)=T(g)\approx T_{c}\approx 0.995$, and $T(d)=T(h)\approx T_{\text{de}}\approx 1.06$. The number of isolated spins reaches a maximum at $T_{\text{ind}}$ and then decreases as temperature increases further, while the number of single spins rises continuously, indicating increasing disorder in the system.
Considering an arbitrary value of $q$, it is observed that for different system sizes $L$, the behavior of $\langle n_{\text{iso}}\rangle/L^{2}$ as a function of temperature remains consistent. That is, for the same value of $q$, the proportion of isolated spins to the system size $L^{2}$ remains invariant across different system sizes. Based on this, it is hypothesized that if the system is divided into several uniform regions, the isolated spins dominate each region. Once an isolated spin appears in a region, the probability of another isolated spin appearing in the same region decreases, and new isolated spins are more likely to appear in previously unoccupied regions. When all regions are occupied by isolated spins, the number of isolated spins reaches its maximum. As the temperature increases, isolated spins tend to reappear in regions where they have already emerged. These newly formed isolated spins merge with the existing ones, forming new clusters, which explains the subsequent decrease in the number of isolated spins.
@@
The extremum of $\langle n_{\text{iso}}\rangle/L^{2}$ is examined for different values of $q$: $\langle n_{\text{iso}}(3)\rangle/L^{2}\approx\langle n_{\text{iso}}(4)\rangle/L^{2}\approx\langle n_{\text{iso}}(6)\rangle/L^{2}\approx\langle n_{\text{iso}}(8)\rangle/L^{2}\approx 0.0215$. This is reasonable because, at $T<T_{c}$, the percolation cluster remains dominant, and in a square lattice system, the size and configuration of the percolation cluster are similar across different values of $q$. Hence, the proportion of isolated spins required to break the percolation cluster is nearly the same. This also explains why the third-order independent phase transition does not depend on the order of the system’s phase transition and is consistently present.
@@
Finally, an explanation is provided for why the temperature associated with the third-order dependent phase transition decreases as $q$ increases. By referring to equation (2) and the SW algorithm, along with Fig. 5, it is noted that when a region is to generate an isolated spin, its spin orientation must differ from that of the percolation cluster. For smaller values of $q$, such as $q=3$, an isolated spin has only two alternative spin orientations, with the third option being reintegration into the percolation cluster. In this case, the probability that a spin becomes isolated is $2/3$, requiring more energy to induce a spin variation. However, for $q=8$, the situation is different. The probability that a spin becomes isolated increases to $7/8$, significantly reducing the energy required. Consequently, the temperature at which isolated spins reach their extremum decreases as $q$ increases.
@@
3.3 Third-order dependent transition in the paramagnetic phase
@@
Fig. 6 illustrates the temperature dependence of the average perimeter $\langle G\rangle$ for the three-state and four-state Potts models. The first row displays the variation of the average perimeter $\langle G\rangle$ with temperature, exhibiting a backbending pattern analogous to that observed in the Ising model. In Fig. 6 (c) and (d), it is observed that $d\langle G\rangle/dT$ exhibits a peak, and as the system size increases, the peak of $d\langle G\rangle/dT$ approaches the system’s critical temperature $T_{c}$, for both $q=3$ and $q=4$. Following this, after $T_{c}$, a minimum is observed in the variation of $d\langle G\rangle/dT$ with temperature, as shown in Fig. 6 (c) and (d). The temperature at which this minimum occurs corresponds to the system’s third-order transition temperature, and these results align with those obtained through microcanonical inflection-point analysis [47].
(a) (a) $T_{c}(3)\approx 0.996$
(b) (b) $T(3)\approx 1.000$
(c) (c) $T_{\text{de}}(3)\approx 1.060$
@@
Figure 7: Clusters identified in typical spin configurations on the 500 × 500 lattice for $q=3$ at three different temperatures: (a)$T\approx 0.996$, is the critical temperature. (b)$T\approx 1.000$, just above the critical temperature, is approaching third dependent transition temperature. (c)$T\approx 1.060$,is the third dependent transition temperature.
@@
By comparing the average perimeter results for the three-state and four-state Potts models in Fig. 6 (a) and (b), we observe a noteworthy phenomenon: the average perimeter $\langle G\rangle$ exhibits a local minimum prior to the critical phase transition and a local maximum subsequent to the transition. The maximum is referred to as the "peak," whereas the minimum is designated as the "valley." The difference between the peak and the valley is termed the "height difference." It is evident that as $q$ increases in the $q$-state Potts model (for $q\geq 3$), this height difference diminishes significantly. This observation reflects the behavior of clusters within the system under varying $q$-state conditions. As $q$ increases, the boundary conditions of the clusters become smoother, resulting in a simplification of the fractal structure. In Fig. 6 (c) and (d), it is observed that $d\langle G\rangle/dT$ exhibits a local minimum. The temperature corresponding to this minimum is identified as the temperature $T_{de}$ associated with the third-order dependent transition. This minimum indicates that, within a temperature range $T$ prior to $T_{de}$, the average perimeter decreases as a concave function with increasing temperature; conversely, after $T_{de}$, it decreases as a convex function. Therefore, $T_{de}$ marks the inflection point at which the concavity of the average perimeter as a function of temperature changes, signifying the location of the third-order dependent transition. To gain a more intuitive understanding of the cluster information within the system, we simulated the clusters in the $q=3$ Potts model at three distinct temperatures: $T_{c}(3)\approx 0.996$, $T(3)\approx 1.000$, and $T_{de}(3)\approx 1.060$, as shown in Fig. 7.
@@
In Fig. 8 (a) and (b), it is observed that the average perimeter $\langle G\rangle$ decreases as the temperature increases. Although the first derivative $d\langle G\rangle/dT$ exhibits a minimum, as shown in Fig. 8 (c) and (d), it is evident that this minimum shifts closer to the system’s critical temperature $T_{c}$ as the system size increases, without indicating a signal of a third-order dependent transition. Previously, the order of the phase transition was determined by analyzing the relationship between $\langle P_{\infty}\rangle$ and $\langle M\rangle$, as shown in Fig. 3. For $q=6$ and $q=8$, the system undergoes a first-order phase transition. Based on this, it is speculated that no third-order dependent transition signals exist in systems undergoing first-order transitions. In contrast, for $q=3$ and $q=4$, where the system exhibits a continuous phase transition, third-order dependent transition signals are present.
To further investigate the suitability of the average perimeter $\langle G\rangle$ as an order parameter, calculations were also performed for the Ising model. Notably, the temperature at which the average perimeter changes is slightly higher than the temperature corresponding to the third-order dependent transition when calculated using the average cluster size. However, the difference between the two temperatures is not significant. This discrepancy could be attributed to the unique two-state nature of the Ising model, in which noise in the canonical ensemble may lead to such variations between the two methods.
4 Conclusion

### v3

@@ III Results and discussion
The following section discusses the results obtained from Swendsen-Wang spin cluster simulations and cluster analysis to elucidate the system behavior associated with the additional transitions in the Potts model, as identified by microcanonical inflection-point analysis.
III.1 Phase transition
@@
Figure 3: The variation of $\langle P_{\infty}\rangle$ with temperature is shown. The black dashed line marks the precise solution of the system’s phase transition temperature $T_{\text{c}}$.As shown in panel (a) and (b), it is observed that at $T_{\text{c}}$, the $q=3,4$ Potts models exhibit characteristics of a continuous phase transition, indicating that the system undergoes a second-order phase transition for $q=3,4$. In contrast, in panel (c) and (d), for $q=6,8$, $\langle P_{\infty}\rangle$ shows a discontinuity at $T_{\text{c}}$, signifying that the system undergoes a first-order phase transition for $q=6,8$.
@@
The primary goal of this paper is to determine the locations of both the third-order independent and dependent transitions in the $q$-state Potts model (for $q\geq 3$) for finite system sizes. The system sizes studied in this work are based on square lattices with edge lengths $L=32,64,128,256$. As derived by Wu Wu (1982), the exact solution for the phase transition temperature $T_{\text{c}}$ of the $q$-state Potts model (for $q\geq 3$) is given by:
$T_{\text{c}}=\frac{1}{\ln(1+\sqrt{q})}$ (6)
@@
Therefore, for $q=3,4,6,8$, the corresponding phase transition temperatures approximate $T_{\text{c}}(3)\approx 0.995$, $T_{\text{c}}(4)\approx 0.910$, $T_{\text{c}}(6)\approx 0.808$, and $T_{\text{c}}(8)\approx 0.745$, respectively.
@@
After identifying all the clusters in the system, various types of clusters emerge, and it is inevitable that the system contains a largest spin cluster. The phase behavior of $\langle P_{\infty}\rangle$ closely mirrors that of $\langle M\rangle$ for $T\leq T_{\text{c}}$, with minor deviations observed for $T>T_{\text{c}}$, while the determination of whether the phase transition is continuous remains largely consistent Kasteleyn and Fortuin (1969). Consequently, analyzing the phase transition behavior of $\langle P_{\infty}\rangle$ facilitates identifying the order of the phase transition in the $q$-state Potts model (for $q\geq 3$) Baxter (2016). In Fig. 3 (a) and (b), it is observed that for $q=3$ and $q=4$, $\langle P_{\infty}\rangle$ changes continuously with temperature, which aligns with the results from Wu Wu (1982) and Baxter Baxter (1973). This confirms that the phase transitions in the $q$-state Potts model for $q=3$ and $q=4$ are critical, i.e., second-order phase transitions. For $q=6$ and $q=8$, $\langle P_{\infty}\rangle$ exhibits discontinuous behavior, consistent with the numerical simulation results reported in the existing literature Baxter (1973); Iglói and Carlon (1999); Hamer (1981); Iglói and Sólyom (1983); Loureiro et al. (2010, 2012), indicating a first-order phase transition when $q=6$ and $q=8$.
@@
Figure 4: The variation of isolated spins (ISO) with temperature for $q=3,4,6,8$ is depicted in panels (a)-(d), respectively. The purple dashed line marks the temperature corresponding to the third-order independent transition, while the black dashed line indicates the precise solution of the system’s phase transition temperature $T_{\text{c}}$. As shown in the figure, the temperatures corresponding to the third-order independent transition are identified as $T_{\text{ind}}(3)=0.964$, $T_{\text{ind}}(4)=0.880$, $T_{\text{ind}}(6)=0.788$, and $T_{\text{ind}}(8)=0.734$.
III.2 Third-order independent transition in the ferromagnetic phase
@@
As shown in Fig. 4 (a)-(d), the number of isolated spins in the $q=3,4,6,8$ Potts models, respectively, varies as a function of temperature. It is observed that the number of isolated spins initially increases with temperature, reaching a peak at $T_{\text{ind}}(3)\approx 0.964$ for the 3-state Potts model, $T_{\text{ind}}(4)\approx 0.880$ for the 4-state Potts model, $T_{\text{ind}}(6)\approx 0.788$ for the 6-state Potts model, and $T_{\text{ind}}(8)\approx 0.734$ for the 8-state Potts model. In Fig. 4, comparing the temperature of the third-order independent transition, $T_{\text{ind}}$, with the phase transition temperature, $T_{\text{c}}$, at the same $q$, the differences are approximately $\Delta T(3)\approx 0.031$, $\Delta T(4)\approx 0.030$, $\Delta T(6)\approx 0.020$, and $\Delta T(8)\approx 0.011$. It is evident that for $q=3$ and $q=4$, the third-order independent transition occurs at a notable distance from the phase transition, with $\Delta T(3)\approx\Delta T(4)$. However, for $q=6$ and $q=8$, $\Delta T$ starts to decrease, and as $q$ increases, $\Delta T$ becomes smaller. Although the temperatures of the third-order independent transition and the system’s phase transition become closer, microcanonical inflection point analysis shows a re-entrant behavior of the inverse temperature as energy increases Wang et al. (2024), meaning that a single temperature corresponds to two energy levels. Near $T=T_{\text{c}}$, the system exhibits a two-phase coexistence. Therefore, despite the proximity of these temperatures, there remains a distinct difference in the system’s states.
@@
Figure 5: A comparison of isolated spins (ISO) and single spins in the $q=3$ Potts model is provided on a lattice with $N=200\times 200$. In the figure, each column compares isolated spins and single spins at the same temperature and configuration. The largest cluster size (LCS) in the system is shaded in gray, while all other clusters containing more than two spins are depicted in white, regardless of their spin orientations. The red points represent spins with $s_{i}=0$, the green points represent $s_{i}=1$, and the blue points represent $s_{i}=2$.
@@
As shown in Fig. 5, it becomes apparent that at low temperatures ($T<T_{\text{c}}$), the largest cluster dominates the system, nearly occupying the entire space, with little variation in its size and configuration. Prior to the phase transition, the percolation cluster continues to dominate, oscillating between similar structures. Disrupting the largest cluster at low temperatures can be achieved by isolated spins with orientations different from those within the largest cluster. During the spin-flip process, the largest cluster has a probability of being flipped, resulting in changes to the required isolated spin states,as shown in Fig. 5 (a)-(b) and (e)-(f). As the temperature increases, the size of the largest cluster decreases. At the phase transition temperature $T_{\text{c}}$, its size can only maintain a span across the entire lattice, but its dominance is much weaker compared to the low-temperature state. As shown in Fig. 5 (c) and (g), with further increase in temperature, the largest cluster begins to collapse rapidly. From Fig. 5 (d) and (h), we can observe that the size of the largest cluster becomes very small, and its size is almost indistinguishable from that of other clusters in the system. In the disordered phase, there is no percolating cluster left in the system, and at this point, studying the influence of the largest cluster size on the system in the disordered phase becomes meaningless. In the paramagnetic phase at high temperatures, the system becomes highly disordered, forming many small clusters and a large number of single spins. Panels (a)-(d) show the variation of isolated spins with temperature, while panels (e)-(f) illustrate the changes in single spin clusters with temperature. The number of isolated spins reaches a maximum at $T_{\text{ind}}$ and then decreases as temperature increases further, while the number of single spins rises continuously, indicating increasing disorder in the system.
For an arbitrary value of $q$, the behavior of $\langle n_{\text{iso}}\rangle/L^{2}$ as a function of temperature remains consistent across different system sizes $L$. This implies that for the same value of $q$, the proportion of isolated spins to the system size $L^{2}$ is invariant across different system sizes. Based on this, combined with Fig. 5, we observe that isolated spins are essentially a special type of single-spin cluster. A single spin appearing in the system can be considered a "mutant" spin. If this mutant spin emerges in an ordered region, it becomes an isolated spin. However, if it appears in a disordered region, it is merely a simple single-spin cluster. At low temperatures, as illustrated in Fig. 5 (a) and (e), isolated spins and single-spin clusters are nearly indistinguishable. As the temperature increases, the number of isolated spins begins to grow, with some merging into new small clusters. In certain cases, isolated spins with different orientations may cluster together, at which point they are no longer considered isolated spins because they disrupt the surrounding region, rendering it disordered. Consequently, they are reclassified as simple single-spin clusters. When the system’s temperature exceeds $T_{\text{ind}}$, the system’s order begins to collapse rapidly. At this stage, ordered regions in the system are quickly disrupted by isolated spins, leading to an increase in the number of small clusters and single-spin clusters, while the number of isolated spins decreases sharply. This observation underscores the necessity of redefining isolated spins. By adopting this redefinition, the physical significance of isolated spins can be better explained, and their behavior before the system’s transition point becomes more pronounced, enabling them to serve as precursors to the system’s transition behavior.
@@
The extremum of $\langle n_{\text{iso}}\rangle/L^{2}$ is examined for different values of $q$: $\langle n_{\text{iso}}(3)\rangle/L^{2}\approx\langle n_{\text{iso}}(4)\rangle/L^{2}\approx\langle n_{\text{iso}}(6)\rangle/L^{2}\approx\langle n_{\text{iso}}(8)\rangle/L^{2}\approx 0.0215$. This is reasonable because, at $T<T_{\text{c}}$, the percolation cluster remains dominant, and in a square lattice system, the size and configuration of the percolation cluster are similar across different values of $q$. Hence, the proportion of isolated spins required to break the percolation cluster is nearly the same. This also explains why the third-order independent transition does not depend on the order of the system’s phase transition and is consistently present.
@@
Finally, an explanation is provided for why the temperature associated with the third-order dependent transition decreases as $q$ increases. By referring to equation (2) and the SW algorithm, along with Fig. 5, it is noted that when a region is to generate an isolated spin, its spin orientation must differ from that of the percolation cluster. For smaller values of $q$, such as $q=3$, an isolated spin has only two alternative spin orientations, with the third option being reintegration into the percolation cluster. In this case, the probability that a spin becomes isolated is $2/3$, requiring more energy to induce a spin variation. However, for $q=8$, the situation is different. The probability that a spin becomes isolated increases to $7/8$, significantly reducing the energy required. Consequently, the temperature at which isolated spins reach their extremum decreases as $q$ increases.
@@
III.3 Third-order dependent transition in the paramagnetic phase
@@
Fig. 6 illustrates the temperature dependence of the average perimeter $\langle G\rangle$ for the three-state and four-state Potts models. The first row displays the variation of the average perimeter $\langle G\rangle$ with temperature, exhibiting a backbending pattern analogous to that observed in the Ising model. In Fig. 6 (c) and (d), it is observed that $d\langle G\rangle/dT$ exhibits a peak, and as the system size increases, the peak of $d\langle G\rangle/dT$ approaches the system’s transition temperature $T_{\text{c}}$, for both $q=3$ and $q=4$. Following this, after $T_{\text{c}}$, a minimum is observed in the variation of $d\langle G\rangle/dT$ with temperature, as shown in Fig. 6 (c) and (d). The temperature at which this minimum occurs corresponds to the system’s third-order transition temperature, and these results align with those obtained through microcanonical inflection-point analysis Wang et al. (2024).
@@
Figure 7: Clusters identified in typical spin configurations on the 500 × 500 lattice for $q=3$ at three different temperatures: (a)$T\approx 0.996$, is the phase transition temperature. (b)$T\approx 1.000$, just above the phase transition temperature, is approaching third dependent transition temperature. (c)$T\approx 1.060$, is the third dependent transition temperature.
@@
By comparing the average perimeter results for the three-state and four-state Potts models in Fig. 6 (a) and (b), we observe a noteworthy phenomenon: the average perimeter $\langle G\rangle$ exhibits a local minimum prior to the phase transition and a local maximum subsequent to the transition. The maximum is referred to as the "peak," whereas the minimum is designated as the "valley." The difference between the peak and the valley is termed the "height difference." It is evident that as $q$ increases in the $q$-state Potts model (for $q\geq 3$), this height difference diminishes significantly. This observation reflects the behavior of clusters within the system under varying $q$-state conditions. As $q$ increases, the boundary conditions of the clusters become smoother, resulting in a simplification of the fractal structure. In Fig. 6 (c) and (d), it is observed that $d\langle G\rangle/dT$ exhibits a local minimum. The temperature corresponding to this minimum is identified as the temperature $T_{de}$ associated with the third-order dependent transition. This minimum indicates that, within a temperature range $T$ prior to $T_{de}$, the average perimeter decreases as a concave function with increasing temperature; conversely, after $T_{de}$, it decreases as a convex function. Therefore, $T_{de}$ marks the inflection point at which the concavity of the average perimeter as a function of temperature changes, signifying the location of the third-order dependent transition. To gain a more intuitive understanding of the cluster information within the system, we simulated the clusters in the $q=3$ Potts model at three distinct temperatures: $T_{\text{c}}(3)\approx 0.996$, $T(3)\approx 1.000$, and $T_{de}(3)\approx 1.060$, as shown in Fig. 7.
@@
In Fig. 8 (a) and (b), it is observed that the average perimeter $\langle G\rangle$ decreases as the temperature increases. Although the first derivative $d\langle G\rangle/dT$ exhibits a minimum, as shown in Fig. 8 (c) and (d), it is evident that this minimum shifts closer to the system’s phase transition temperature $T_{\text{c}}$ as the system size increases, without indicating a signal of a third-order dependent transition. Previously, the order of the phase transition was determined by analyzing the relationship between $\langle P_{\infty}\rangle$ and $\langle M\rangle$, as shown in Fig. 3. For $q=6$ and $q=8$, the system undergoes a first-order phase transition. Based on this, it is speculated that no third-order dependent transition signals exist in systems undergoing first-order transitions. In contrast, for $q=3$ and $q=4$, where the system exhibits a continuous phase transition, third-order dependent transition signals are present.
IV Conclusion